Consider the quadratic equation y=8x^2-2x. Complete the square to express the quadratic in standard form y=a(x-h)^2+k.
I don't understand how to do this problem. The "8" is stumping me.
Thanks for your help

Answer by John10(297)

Consider the quadratic equation y=8x^2-2x. Complete the square to express the quadratic in standard form y=a(x-h)^2+k.
-------------------------------------------------------
For this problem,
we will need to factor first
y = 8(x^2 - x/4)= 8(x^2 - 2(1/8)(x) + 1/64)- 1/64 = 8(x - 1/8)^2 - 1/64
Hope it helps you!
John10(john100185@yahoo.com)